OpenStudy (anonymous):

For what values of m does the graph of y = 3x2 + 7x + m have two x-intercepts? a) m>12/49 b) m<12/49 c) m<49/12 d) m> 49/12

OpenStudy (anonymous):

Do you have any pictures for this question?

OpenStudy (anonymous):

Okay, perfect :)

OpenStudy (anonymous):

Hold on, I'm working it out on paper.

OpenStudy (anonymous):

alright c:

OpenStudy (anonymous):

Alrighty. The quadratic formula (a part we call the "discriminant") is defined by the variables that are inside the square root, and is denotated by the delta

OpenStudy (anonymous):

Δ=b2−4ac

OpenStudy (anonymous):

Whenever we solve a quadratic equation that is complete and we analyze the discriminant,

OpenStudy (anonymous):

we get 3 scenarios: if→Δ>0=>∃x1,x2/ax2+bx+c=0

OpenStudy (anonymous):

This just means: "if the discriminant is greater than zero, there will exist two x-intercepts" And for the second scenario: if→Δ=0→∃xo/ax2+bx+c=0

OpenStudy (anonymous):

This means that if the discriminant is equal to zero, there will be one and only one x-intercept And for the last scenario: (if→Δ<0→∃x∉R/ax2+bx+c=0) This means that if our discriminant is less than zero, there will be no x-intercepts So, if we take your excercise and analyze the the discriminant (which is 3x2+7x+m=y) we can find the values that satisfy y=0 3x2+7x+m=0 than we can analyze the discriminant: Δ=72−4(3)(m) we are only interested in the values that make the discriminant equal zero 72−4(3)(m)=0 so all you have to do is solve for "m"

OpenStudy (anonymous):

im very confused

OpenStudy (anonymous):

Oh no :( I'll simplify this to find the intercepts of the graph, you have to solve for "m"

OpenStudy (anonymous):

D is your answer